to the nearest hundredth, find the surface area of a cone whose slant height is 9 centimeters and that has a base with radius 6 centimeters.
Answer by checkley71(8403)

TSA=1/2PS+BASE
TSA=1/2(PI*D)S+PI*R^2
TSA=.5*3.14*12*9+3.14*6^2
TSA=169.56+113.04
TSA=282.6 ANSWER.